Use the discriminant to determine how many real number solutions the quadratic equation -4x^2 + 3x - 28 =0 has.
Answer by stanbon(75887)

Use the discriminant to determine how many real number solutions the quadratic equation -4x^2 + 3x - 28 =0 has.
-------
a = -4 ; b = 3 ; c = -28
----
b^2 - 4ac = 9 - 4*-4*-28 < 0
====
Since the discriminant is < 0, the quadratic
has 2 unequal complex roots.
=========================================
Cheers,
Stan H.